# Two-sample comparison through additive tree models for density ratios

### v1

@@ Abstract
The ratio of two densities characterizes their differences. We consider learning the density ratio given i.i.d. observations from each of the two distributions. We propose additive tree models for the density ratio along with efficient algorithms for training these models using a new loss function called the balancing loss. With this loss, additive tree models for the density ratio can be trained using algorithms original designed for supervised learning. Specifically, they can be trained from both an optimization perspective that parallels tree boosting and from a (generalized) Bayesian perspective that parallels Bayesian additive regression trees (BART). For the former, we present two boosting algorithms—one based on forward-stagewise fitting and the other based on gradient boosting, both of which produce a point estimate for the density ratio function. For the latter, we show that due to the loss function’s resemblance to an exponential family kernel, the new loss can serve as a pseudo-likelihood for which conjugate priors exist, thereby enabling effective generalized Bayesian inference on the density ratio using backfitting samplers designed for BART. The resulting uncertainty quantification on the inferred density ratio is critical for applications involving high-dimensional and complex distributions in which uncertainty given limited data can often be substantial. We provide insights on the balancing loss through its close connection to the exponential loss in binary classification and to the variational form of $f$-divergence, in particular that of the squared Hellinger distance. Our numerical experiments demonstrate the accuracy and computational efficiency of the proposed approach while providing unique capabilities in uncertainty quantification. We demonstrate the application of our method in a case study involving assessing the quality of generative models for microbiome compositional data.
@@ 1 Introduction
A vast range of applications involve comparing two groups of observations to assess their differences. Classical formulation of two-sample comparison often involves parametric or semiparametric assumptions on the underlying distributions and focuses on testing the null hypothesis of no difference with or without a given alternative in mind. However, this hypothesis testing perspective is often inadequate in most modern applications of two-sample comparison. The practitioners nowadays typically also desire to learn the specific nature of the underlying difference. This problem lies at the core of numerous other modern statistical analyses and machine learning objectives.
@@
For example, large-scale biomedical studies involving comparing genetic measurements and various biomarkers of patients and controls, there is often no question that some types of two-sample differences do exist, and with a sufficiently large sample size, the null of no difference at all will be rejected. It is more helpful if one can specifically pinpoint what such differences are to help understand their biological meaning. Another timely application is the assessment of generative models. The key to assessing the quality of different models and algorithms for generating synthetic samples that imitate the real sampling distribution of the training data lies in deciphering the distributional difference between the real sampling distribution and that induced by the generative models, which is essentially a two-sample comparison problem. Some generative models, such as generative adversarial networks (GANs) (Goodfellow et al., 2014), in essence iteratively apply two-sample comparison and use the estimated misfit between the two distributions to sequentially improve the generative model. Two-sample comparison is also a quintessential problem in causal inference, in which consistent estimation of treatment effects often relies on correctly adjusting for the difference in the covariate distributions between the treated and the control groups.
@@
Our focus here is on two-sample comparison in a nonparametric fashion using the density ratio function as a functional “statistic” for characterizing two-sample differences. The density ratio fully summarizes how two distributions differ across the sample space, and thus allows one to not only carry out hypothesis testing but to identify the nature of the difference as well. If, for example, the estimated density ratio “significantly” deviates away from 1 on parts of the sample space, it will allow us to draw statistical conclusions regarding whether, where, and how the two distributions differ. At the same time, the notion of “significant” deviation will require some form of uncertainty quantification on the inferred density ratio.
@@
There is a general misconception that nonparametric density ratio learning is “too hard” for complex, high-dimensional data because density estimation itself in such settings is already a hard enough problem in multivariate. Why then would one attempt an even harder problem of learning the ratio of two densities? Though it might appear counter-intuitive, we argue that density ratio estimation (DRE) is in fact a much easier problem than density estimation in most modern applications. A density by definition is always a density ratio, with the denominator being the corresponding base measure against which the density is computed. The typical base measure, for example the Lebesgue measure or the uniform density in continuous cases, is generally chosen out of mathematical convenience and the lack of knowledge that takes no account of the dataset at hand. It generally does not resemble any sampling distribution in high-dimensional problems. When good reliable knowledge of the sampling distribution does exist, one would be in the classical setting of a “goodness-of-fit” problem. In contrast, in most applications of two-sample comparison, the two distributions are often quite similar to each other and can serve as a “baseline of reference” for each other. As such, the density ratio between these two samples then is often much simpler than the underlying densities (i.e., their ratios with respect to the Lebesgue measure).
@@
This is the fundamental reason behind a well-known observation in the DRE literature, that though estimating a density ratio can be done by first separately estimating the two densities and then computing their ratios based on the density estimates, it is usually more effective to adopt a method directly aimed at the density ratio, especially when the difference exists only in a small subset of the sample space or in a marginal distribution. Several authors have adopted this strategy. Some propose kernel-based methods (e.g. Sugiyama et al., 2007b; Kanamori et al., 2009; Hido et al., 2011; Liu et al., 2013; Yamada et al., 2013), and various kinds of other methods are proposed, such as an estimation via classification (Rhodes et al., 2020; Choi et al., 2021, 2022) and a deep neural network method (Kato and Teshima, 2021).
@@
Beyond the stand-alone two-sample comparison problem which is our focus herein, DRE has also been applied as a component of various types of learning methods in which two-sample comparison arises as a sub-problem directly or indirectly. A far-from-exhaustive list includes an estimation of the mutual information (Suzuki et al., 2009a, b), conditional density estimation (Sugiyama et al., 2010), covariate shift correlation (Shimodaira, 2000; Sugiyama et al., 2007a; Reddi et al., 2015; Stojanov et al., 2019), outlier detection (Hido et al., 2011; Kato and Teshima, 2021), causal inference (Uehara et al., 2020), and change point detection (Kawahara and Sugiyama, 2009; Liu et al., 2013; du Plessis et al., 2015). Additionally, estimating density ratios is utilized in Markov Chain Monte Carlo (MCMC) for intractable likelihoods (Kaji and Ročková, 2023).
@@
We list several desiderata in learning the density ratio and discuss how we aim to achieve them. First, the model or function class representing the density ratio should be flexible enough to effectively approximate complex forms of the underlying truth. It should scale well in both sample size and in dimensionality. To these ends, we propose a class of density ratio learners based on additive tree boosting (see e.g. Friedman et al., 2000; Hastie et al., 2009). Additive tree ensembles are a natural candidate for the density ratio problem, because of their success in classification problems. In fact, in the machine learning literature, a widely used technique to learn the density ratio between two samples is the so-called “density-ratio trick” (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022), which involves first fitting a binary classifier to the two samples, and then, based on the estimated posterior odds of the sample label, compute the density ratio. Additive tree boosting such as AdaBoost is among the most effective for binary classification on the so-called “tabular data”. We show that DRE does not have to involve the indirect route of inverting a binary classifier based on additive tree ensembles, but can directly be learned from the latter. As we will show, the resulting estimate will often be substantially more accurate than what the density ratio trick will produce, especially when the sample size is unbalanced or the two-sample difference is highly local; in both of these scenarios, losses designed for binary classification are a poor proxy for learning the density ratio.
@@
We introduce two tree boosting algorithms for estimating the density ratio. We use an additive tree ensemble as the function class for approximating the density ratio. by adding tree-based weak learners. The first algorithm is based on forward-stagewise fitting (Hastie et al., 2009) and it operationally resembles Adaboost. In each iteration the observations are weighted according to the discrepancies between the current estimate and the actual responses, so that in each step we can fit a new weak learner focusing on observations with large discrepancies. This algorithm sequentially and greedily minimizes the balancing loss through fitting a single-tree piecewise constant approximation to the density ratio of the weighted samples. The second algorithm is under the general paradigm of gradient boosting (Friedman, 2001).
@@
These algorithms inherit important advantages of supervised tree boosting — they can be implemented with a small computation cost, and at the same time achieve good numerical performance, especially when the difference lies in a marginal distribution of a subset of the variables. One can incorporate regularization to avoid overfitting effectively by making the contribution of each base learner weak with a small learning rate, following the standard approach in supervised boosting (Hastie et al., 2009; Friedman, 2001).
@@
The second desideratum is to attain uncertainty quantification for the estimated density ratio, thereby allowing statistical inference to proceed. This aspect has mostly been ignored in the existing literature on DRE. One exception was Bayesian nonparametric models for the DR using a single tree model (Ma and Wong, 2011; Soriano and Ma, 2017). While they provide means to examine posterior uncertainty in a fully Bayesian manner, piecewise-constant functions represented by single trees are very restrictive and therefore inadequate for approximating complex density ratios. Additive trees substantially enrich the class of functions, but they do not allow a simple likelihood formulation for modeling density ratios, and consequently, fully Bayesian uncertainty quantification is challenging. To overcome this, we propose a generalized Bayesian approach Bissiri et al. (2016) using the balancing loss to construct a pseudo-likelihood for this model class that has shown excellent performance in supervised problems, as realized by Bayesian additive and regression trees (BART) (Chipman et al., 2010). Interestingly, we show that due to the resemblance of the balancing loss to an exponential family kernel, there exists a natural conjugate prior for the model parameters under the pseudo-likelihood which allows standard sampling strategies for BART models (Chipman et al., 2010) to be directly adopted for sampling the Gibbs posterior on the density ratio.
@@
The remainder of the paper is organized as follows. In Section 2 we start by introducing the balancing loss for learning the density ratio and draw its connection to the variational form of squared Hellinger distance and to the exponential loss for binary classification. We then introduce an additive tree model on the (log) density ratio along with two boosting algorithms for fitting this model based on forward stage-wise fitting and gradient boosting, respectively. In Section 3, we introduce a generalized Bayesian inference framework that uses the balancing loss as a (log) pseudo-likelihood and show one can accomplish Bayesian backfitting sampling on the model using standard algorithms for BART through exploiting the analogy of the pseudo-likelihood to an exponential family. In Section 4, we evaluate the performance of the proposed methods in comparison with the existing approaches in lower and higher-dimensional simulation settings followed by a case study involving assessing and comparing multiple generative models for microbiome compositions in Section 5. We give a few concluding remarks in Section 6.
@@ 2.1 The balancing loss for density ratio learning
When $B=\Omega$, the whole sample space, this corresponds to the affinity between $p$ and $q$. The first equation implies that if we weight the two samples respectively with $r^{-1/2}$ and $r^{1/2}$, their probability over any set $B$ becomes equal. In this sense, we say the two samples are “balanced”. Motivated by this observation, we introduce the notation $w=r^{1/2}$ to denote the balancing function, which is intended to approximate the square root of the density ratio $\sqrt{p/q}$. We then consider the following balancing loss function:
@@
In practice, we observe i.i.d. draws from $p$ and $q$:
@@
The finite-sample version of the balancing loss is defined by
@@
where $n=n^{(0)}+n^{(1)}$ indicates the total number of observations.
@@ 2.2 Connections to existing concepts in statistical learning
While the design of the balancing loss may appear ad hoc at first, next we show that it can be naturally motivated from two well-established perspectives in machine learning.
@@ 2.2.1 Connection 1: the “density ratio trick” based on binary classification
The first perspective is related to a common technique widely adopted in the machine learning community for estimating the density, often referred to as the “density ratio trick”, which involves turning the problem into a supervised, binary classification on the two group labels (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022). It involves inverting a binary classifier to compute the density ratio. We show that the balancing loss arises naturally in this context as a variant of the exponential loss widely used in this context, but only the balancing loss targets the density ratio directly and thus is more robust than the density ratio trick when the two sample sizes are unbalanced and/or the two-sample differences exist locally in a small subset of the sample space. These will also be validated experimentally later in our numerical examples where a direct comparison to the density ratio trick will be provided.
The balancing loss can be seen as a modification of the exponential loss, which is used for classification with the Adaboost, described in Friedman et al. (2000). To see this, we suppose that in generating the two samples, one first randomly decides which sample to draw the observation from. That is, one first draws a random variable $y$ that takes 1 and -1 with a 50-50 chance. Given $y$ one generates the corresponding sample $x$ from $P$ if $y=1$ and from $Q$ if y=1y=-1, respectively. Then it is straightforward to show that
@@
where $F=\log w$ and the expectation in the latter is taken with respect to the joint distribution of $x$ and $y$. What we have on the right-hand side is exactly the exponential loss (Friedman et al., 2000; Hastie et al., 2009).
We note that the exponential loss used in AdaBoost in fact integrates $y$ with respect to its marginal distribution from the data. While the population minimizer for the AdaBoost is the log odds between the two samples (Friedman et al., 2000), which is proportional to the density ratio, one may wonder then why the original AdaBoost cannot suffice for learning density ratios since one can simply multiply a constant (i.e., the ratio of the proportion of the two samples) to get an estimate of the density ratio. The main reason is that by weighting the loss using the marginal proportion of two samples, vanilla AdaBoost prioritizes classification accuracy over density ratio estimation; this can be most clearly demonstrated when there are parts of the sample space in which one of the samples is much more abundant than the other, which can occur when the sample sizes are very different or when the two sample densities have different supports or tail behaviors. Such cases are in fact quite common in multivariate applications, and numerical evaluations will be provided in Section 4.
@@
The design of the balancing loss can also be motivated from the perspective of estimating the $f$-divergence between $p$ and $q$. In particular, in this section we show that the balancing loss turns out to be exactly equivalent to the variational form of one particular $f$-divergence, namely the squared Hellinger distance.
@@
which is clearly equivalent to minimizing the balancing loss. In other words, learning the density ratio by minimizing the balancing loss is essentially estimating the squared Hellinger distance using its variational form. Based on the known minimizer of the variational form, the minimizer of the balancing loss is $w^{*}=\sqrt{p/q}$. Hence, learning the density ratio can be achieved through minimizing the balancing loss.
Beyond these two additional motivations of the balancing loss, we will see later that the symmetry of the balancing loss, in $w$ vs $w^{-1}$ and $p$ vs $q$ will become handy when designing a generalized Bayesian framework to carry out inference on additive tree models.
2.3 Choice of the function class: additive tree models
Given the close relationship between density ratios and posterior odds, the latter of which can be estimated using supervised tree boosting with additive tree models, it is natural to consider the class of additive trees as a promising candidate for approximating the density ratio. We thus consider from now on fitting additive tree models to the density ratio under the balancing loss.
We begin with a non-Bayesian approach that aims to find a (functional) point estimate of the density ratio by optimizing the finite-sample version of the balancing loss. The resulting training algorithms are very similar to supervised tree boosting. Before introducing the algorithms, we describe the additive tree model. We consider a model for the log-balancing weight $\log w$ in the form

### v2

@@ Abstract
The ratio of two densities characterizes their differences. We consider the two-sample comparison problem through estimating the density ratio given i.i.d. observations from each distribution. We propose additive tree models for the density ratio along with efficient algorithms for training these models using a new loss function called the balancing loss. With this loss, additive tree models for the density ratio can be trained using several algorithms originally designed for supervised learning. First, they can be trained from an optimization perspective through boosting algorithms including forward-stagewise fitting and gradient boosting. Moreover, due to the balancing loss function’s resemblance to an exponential family kernel, it can serve as a pseudo-likelihood for which conjugate priors exist, thereby enabling effective generalized Bayesian inference on the density ratio using backfitting samplers designed for Bayesian additive regression trees (BART). This generalized Bayesian strategy enables uncertainty quantification on the inferred density ratio, which is critical for applications involving high-dimensional and complex distributions, as in such problems uncertainty given limited data can often be substantial. We provide insights on the balancing loss through its close connection to the exponential loss in binary classification and to the variational form of $f$-divergence, in particular that of the squared Hellinger distance. Our numerical experiments demonstrate the accuracy and computational efficiency of the proposed approach while providing unique capabilities in uncertainty quantification. We demonstrate the application of our method in a case study involving assessing the quality of generative models for microbiome compositional data.
@@ 1 Introduction
A vast range of applications involve comparing two groups of observations to assess their differences. The classical formulation of the two-sample comparison often involves parametric or semiparametric assumptions on the underlying distributions and focuses on testing the null hypothesis of no difference with or without a given alternative in mind. However, this hypothesis testing perspective is often inadequate for most modern applications of two-sample comparison. Practitioners today typically also desire to learn the specific nature of the underlying difference. This problem lies at the core of numerous other modern statistical analyses and machine learning objectives.
@@
For example, in large-scale biomedical studies involving the comparison of genetic measurements and various biomarkers between patients and controls, there is often no question that some types of two-sample differences do exist, and with a sufficiently large sample size, the null of no difference at all will be rejected. It is more helpful to specifically pinpoint what such differences are in order to help understand their biological meaning. Another timely application is the assessment of generative models. The key to assessing the quality of different models and algorithms for generating synthetic samples that imitate the real sampling distribution of the training data lies in deciphering the distributional difference between the real sampling distribution and that induced by the generative models, which is essentially a two-sample comparison problem. Some generative models, such as generative adversarial networks (GANs) (Goodfellow et al., 2014), essentially iteratively apply two-sample comparison and use the estimated misfit between the two distributions to sequentially improve the generative model. Two-sample comparison is also a quintessential problem in causal inference, in which consistent estimation of treatment effects often relies on correctly adjusting for the difference in the covariate distributions between the treated and the control groups.
@@
Our focus here is on carrying out two-sample comparison in a nonparametric fashion through estimation rather than hypothesis testing, using the density ratio function as the functional “statistic” of interest for characterizing two-sample differences. The density ratio fully summarizes how two distributions differ across the sample space, and thus allows one not only to carry out hypothesis testing but also to identify the nature of the difference as well. If, for example, the estimated density ratio deviates “significantly” away from 1 on parts of the sample space, it allows us to draw statistical conclusions regarding whether, where, and how the two distributions differ. At the same time, the notion of “significant” deviation will require some form of uncertainty quantification for the inferred density ratio, which will become available under our formulation.
@@
One may reasonably raise the concern that nonparametric density ratio learning is simply “too hard” for complex, high-dimensional data because density estimation itself in such settings is already a challenging problem. Why then would one attempt an even harder problem of learning the ratio of two densities? Though it might appear counter-intuitive, we argue that density ratio estimation (DRE) is in fact a much easier problem than density estimation in most modern applications. A density by definition is always a density ratio, with the denominator being the corresponding base measure against which the density is computed. The typical base measure, for example the Lebesgue measure or the uniform density in continuous cases, is generally chosen out of mathematical convenience and lack of prior knowledge, without taking any account of the dataset at hand (other than that the observations are continuous). This base measure generally does not resemble any sampling distribution in high-dimensional or otherwise complex problems. When good reliable knowledge of the sampling distribution does exist, one would be in the much simpler classical setting of a “goodness-of-fit” problem. In contrast, in most applications of two-sample comparison, the two distributions are often quite similar to each other and can serve as a “baseline of reference” for each other. As such, the density ratio between these two samples then is often much simpler than the underlying densities (i.e., their ratios with respect to the Lebesgue measure).
@@
We believe this explains a well-known observation in the DRE literature, that though estimating a density ratio can be done by first separately estimating the two densities and then computing their ratios based on the density estimates, it is usually more effective to adopt a method directly aimed at the density ratio, especially when the difference exists only in a small subset of the sample space or in a subset of the dimensions. Several authors have adopted this perspective and proposed methods directly targeting the density ratio. Some approaches include kernel-based methods (e.g. Sugiyama et al., 2007b; Kanamori et al., 2009; Hido et al., 2011; Liu et al., 2013; Yamada et al., 2013) and more recently neural network-based methods (Kato and Teshima, 2021).
@@
Beyond the stand-alone two-sample comparison problem which is our focus herein, DRE has also been applied as a component in various learning methods in which two-sample comparison arises as a sub-problem directly or indirectly. A far-from-exhaustive list includes the estimation of the mutual information (Suzuki et al., 2009a, b), conditional density estimation (Sugiyama et al., 2010), covariate shift correction (Shimodaira, 2000; Sugiyama et al., 2007a; Reddi et al., 2015; Stojanov et al., 2019), outlier detection (Hido et al., 2011; Kato and Teshima, 2021), causal inference (Uehara et al., 2020), and change point detection (Kawahara and Sugiyama, 2009; Liu et al., 2013; du Plessis et al., 2015). Additionally, estimating density ratios has been utilized in Markov Chain Monte Carlo (MCMC) for intractable likelihoods (Kaji and Ročková, 2023).
@@
We list several desiderata for learning the density ratio and discuss how we aim to achieve them. First, the model or function class representing the density ratio should be flexible enough to effectively approximate complex forms of the underlying truth. It should scale well with both sample size and in dimensionality. To these ends, we propose a class of density ratio learners based on additive tree ensembles (see e.g., Friedman et al., 2000; Hastie et al., 2009). Additive tree ensembles are a natural candidate for the density ratio problem, because of their success in classification problems. In fact, in the machine learning literature, a widely used technique to learn the density ratio between two samples is the so-called “density-ratio trick” (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022), which involves first fitting a binary classifier to the two samples, and then, based on the estimated posterior odds of the sample label, computing the density ratio. Additive tree boosting such as AdaBoost is among the most effective methods for binary classification.
In contrast to this popular perspective, we hold that DRE should not and does not need to involve the indirect route of inverting a binary classifier based on additive tree ensembles, but can instead be directly learned from the latter. As we will show, the resulting estimate are often substantially more accurate than what the density -ratio trick produces, especially when the sample size is unbalanced and/or the two-sample difference is local. In these scenarios, losses designed for binary classification are poor proxies for learning the density ratio.
@@
In terms of computational strategies, we introduce two boosting algorithms for estimating the additive tree model on density ratios. We use an additive tree ensemble as the function class for approximating the density ratio, by adding tree-based weak learners. The first algorithm is based on forward-stagewise fitting (Hastie et al., 2009) and it operationally resembles AdaBoost. In each iteration, the observations are weighted according to the discrepancies between the current estimate and the actual responses, so that in each step we can fit a new weak learner that focuses on observations with large discrepancies. This algorithm sequentially and greedily minimizes the balancing loss through fitting a single-tree piecewise constant approximation to the density ratio of the weighted samples. The second algorithm follows the general paradigm of gradient boosting (Friedman, 2001).
@@
These algorithms inherit important advantages of supervised tree boosting — they can be implemented at a small computation cost, and at the same time achieve good numerical performance, especially when the difference lies in a marginal distribution of a subset of the variables. One can incorporate regularization to avoid overfitting effectively by making the contribution of each base learner weak through a small learning rate, following the standard approach in supervised boosting (Hastie et al., 2009; Friedman, 2001).
@@
The second desideratum is to attain uncertainty quantification for the estimated density ratio, thereby allowing statistical inference to proceed. This aspect has mostly been ignored in the existing literature on DRE. One exception is Bayesian nonparametric models for the density ratio using a single tree model (Ma and Wong, 2011; Soriano and Ma, 2017). While they provide a means to examine posterior uncertainty in a fully Bayesian manner, piecewise-constant functions represented by single trees of controlled complexity are very restrictive and therefore inadequate for approximating complex density ratios. Additive trees substantially enrich the class of functions, but they do not allow a simple likelihood formulation for modeling density ratios, and consequently, fully Bayesian uncertainty quantification for this function class is challenging. To overcome this, we propose a generalized Bayesian approach (Bissiri et al., 2016) using the balancing loss to construct a pseudo-likelihood for this model class. Interestingly, we show that, due to the resemblance of the balancing loss to an exponential family kernel, there exists a natural conjugate prior for the model parameters under the pseudo-likelihood, which allows standard sampling strategies for BART models (Chipman et al., 2010) to be directly adopted for sampling the Gibbs posterior on the density ratio.
@@
The remainder of the paper is organized as follows. In Section 2 we start by introducing the balancing loss for learning the density ratio and highlight its connection to the variational form of the squared Hellinger distance and to the exponential loss for binary classification. We then introduce an additive tree model for the (log) density ratio along with two boosting algorithms for fitting this model based on forward stage-wise fitting and gradient boosting, respectively. In Section 3, we introduce a generalized Bayesian inference framework that uses the balancing loss as a (log) pseudo-likelihood and show that one can accomplish Bayesian backfitting sampling on the model using standard algorithms for BART by exploiting the analogy of the pseudo-likelihood to an exponential family. In Section 4, we evaluate the performance of the proposed methods in comparison with the existing approaches in a few simulation settings, followed by a case study that assesses and comparing multiple generative models for microbiome compositions in Section 5. We give concluding remarks in Section 6.
@@ 2.1 The balancing loss for density ratio learning
When $B=\Omega$, the whole sample space, this corresponds to the affinity between $p$ and $q$. The first equation implies that if we weight the two samples respectively with $r^{-1/2}$ and $r^{1/2}$, their probability over any measurable set $B$ becomes equal. In this sense, we say that the two samples are “balanced”. Motivated by this observation, we introduce the notation $w=r^{1/2}$ to denote the balancing function, which is intended to approximate the square root of the density ratio $\sqrt{p/q}$. We then consider the following balancing loss function:
@@
In practice, suppose we observe i.i.d. draws from $p$ and $q$:
@@
Then the finite-sample version of the balancing loss is defined by
@@
where $n=n_{0}+n_{1}$ indicates the total number of observations.
@@ 2.2 Connections to existing concepts in statistical learning
While the design of the balancing loss may appear ad hoc at first, we next show that it can be naturally motivated from two well-established perspectives in machine learning.
@@ 2.2.1 Connection 1: the “density ratio trick” based on binary classification
The first perspective is related to a common technique widely adopted in the machine learning community for estimating density ratios, often referred to as the “density-ratio trick”, which involves turning the problem into a supervised, binary classification task on the two group labels (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022). It involves inverting a binary classifier to compute the density ratio. We show that the balancing loss arises naturally in this context as a variant of the exponential loss widely used in this context, but only the balancing loss targets the density ratio directly and thus is more robust than the density ratio trick against scenarios such as when the two sample sizes are unbalanced and/or the two-sample differences exist locally in a small subset of the sample space. These will also be validated experimentally later in our numerical examples where a direct comparison with the density ratio trick will be provided.
The balancing loss can be seen as a modification of the exponential loss, which is used for classification with the AdaBoost, described in Friedman et al. (2000). To see this, we suppose that, in generating the two samples, one first randomly decides which sample to draw the observation from with equal probability. That is, one first draws a random variable $y$ that takes 1 and -1 with a 50-50 chance. Given $y$ one generates the corresponding sample $x$ from $P$ if $y=1$ and from $Q$ if y=1y=-1, respectively. Then it is straightforward to show that
@@
where $F=\log w$, and the expectation in the latter is taken with respect to the joint distribution of $x$ and $y$. What we have on the right-hand side is exactly the exponential loss (Friedman et al., 2000; Hastie et al., 2009).
We note that the exponential loss used in AdaBoost in fact integrates $y$ with respect to its marginal distribution from the data rather than using a 50-50 split. Because the population minimizer for the AdaBoost is the log posterior odds between the two samples (Friedman et al., 2000), which is the product of the density ratio and the prior odds, one may wonder then why the original AdaBoost cannot suffice for learning density ratios since one can simply multiply by a constant (i.e., the ratio of the proportion of the two samples) to get an estimate of the density ratio. The main reason is that, by weighting the loss using the marginal proportion of the two samples, vanilla AdaBoost prioritizes classification accuracy over density ratio estimation; this can be most clearly demonstrated when there are parts of the sample space in which one of the samples is much more abundant than the other, which can occur when the sample sizes are very different or when the two sample densities have different supports or tail behaviors. Such cases are in fact quite common in multivariate applications, and numerical evaluations will be provided in Section 4.
@@
The design of the balancing loss can also be motivated from the perspective of estimating the an $f$-divergence between $p$ and $q$. In particular, in this section we show that the balancing loss turns out to be exactly equivalent to the variational form of one particular $f$-divergence, namely the squared Hellinger distance.
@@
which is clearly equivalent to minimizing the balancing loss. In other words, learning the density ratio by minimizing the balancing loss is essentially estimating the squared Hellinger distance using its variational form. Based on the known minimizer of the variational form, the minimizer of the balancing loss is $w^{*}=\sqrt{p/q}$. Hence, learning the density ratio can be achieved by minimizing the balancing loss.
Beyond these two additional motivations for the balancing loss, we will see later that the symmetry of the balancing loss, in $w$ vs $w^{-1}$ and $p$ vs $q$, will become handy when designing a generalized Bayesian framework to carry out inference on additive tree models.
2.3 Additive tree models and boosting
Given the close relationship between density ratios and posterior odds, the latter of which can be estimated using supervised tree boosting with additive tree models, it is natural to consider the class of additive tree models as a promising candidate for approximating the density ratio. We thus consider from now on fitting additive tree models to the density ratio under the balancing loss.
We begin with a non-Bayesian approach that aims to find a (functional) point estimate of the density ratio by optimizing the finite-sample version of the balancing loss. The resulting training algorithms are very similar to those used in supervised tree boosting. Before introducing the algorithms, we describe the additive tree model. We consider a model for the log-balancing weight $\log w$ in the form